OpenStudy (lxelle):

By expressing 8 sinx - 6 cosx in the form R sin(x - a), solve the equation 8 sinx - 6 cosx = 7, for 0 <= x <= 360

OpenStudy (anonymous):

\[\sqrt{6^2+8^2} = 10\]

OpenStudy (anonymous):

divide the given equation by 10

OpenStudy (lxelle):

i got 10 sin(x - 6)

OpenStudy (anonymous):

\[10 \times (\frac {8 sinx}{10} - \frac{6 cosx}{10}) = 10 \times(sinx \times Cos\alpha- cos x \times sin\alpha)\]here \[cos\alpha = 8/10\]\[sin\alpha = 6/10\]you'd get \[10\times (\frac {8 sinx}{10} - \frac{6 cosx}{10})= 10 sin(x-\alpha)\]

OpenStudy (anonymous):

\[\alpha = sin^{-1}(\frac 6 {10})\]

OpenStudy (lxelle):

uh, i think youre doing it wrong.

OpenStudy (anonymous):